OpenStudy (muscrat123):

Molly's Custom Kitchen Supplies sells homemade forks and spoons. It costs the store $1.70 to buy the supplies to make a fork and $1.30 to buy the supplies to make a spoon. The store sells forks for $5.60 and spoons for 5.40. Last April Molly's Custom Kitchen Supplies spent $37.90 on materials for forks and spoons. They sold the finished products for a total of $147.20. How many forks and how many spoons did they make last April?

OpenStudy (muscrat123):

@Hero

OpenStudy (compassionate):

This is asking for a system of equations, am I right? Are you familiar with system of equations?

OpenStudy (muscrat123):

yea and yea. i know how to do this stuff, its just this one problem

OpenStudy (compassionate):

"Molly's Custom Kitchen Supplies sells homemade forks and spoons. It costs the store $1.70 to buy the supplies to make a fork and $1.30 to buy the supplies to make a spoon. The store sells forks for $5.60 and spoons for 5.40. Last April Molly's Custom Kitchen Supplies spent $37.90 on materials for forks and spoons. They sold the finished products for a total of $147.20. How many forks and how many spoons did they make last April?" So it costs the store: 1.70 and 1.30 f = 1.70 s = 1.30 f + s = 37.90 1.70f + 1.30s = 37.90 In theory, if you solve the system, you can plug each term in and found out how many forks and spoons they bought, then divide it by the sell price

OpenStudy (muscrat123):

i have that

OpenStudy (compassionate):

Would you like to clarify which part you're having trouble with so I may be more direct?

OpenStudy (muscrat123):

well i have this: 1.70f + 1.30s = 37.90 5.60f + 5.40s = 147.20 i dont know where to go from there

OpenStudy (muscrat123):

my teacher said that part was right

hero (hero):

I agree with @muscrat123 setup

OpenStudy (compassionate):

Let's see.... Cost 1.70f + 1.30s = 37.90 Revenue 5.60x + 5.40y = 147.20 So now we have two equations, and we can solve either by either a system of elimination or substitution. It looks like elimination would be easier her, as substitution often requires a variable to already be found 1.70f + 1.30s = 37.90 5.60 + 5.40 = 147.20 If I can multiply our first equation by 5.6 and the second by 1.7, I can cancel some variables 9.52f + 7.28s = 212.24 9.52f+ 9.18s = 250.24 Now I can subtract the first equation from the second. Now that I've set it up, do you think you can show me some work?

OpenStudy (muscrat123):

sure give me like 10 mins is that ok? i have to do something else first.

OpenStudy (compassionate):

I'll here. Just remember, time waits for no one.

OpenStudy (muscrat123):

what does i'll here mean?

OpenStudy (compassionate):

It means my keyboard is messed up. I'll be here.

OpenStudy (muscrat123):

oh ok

OpenStudy (muscrat123):

i got 7 forks and 20 spoons f = 7 s = 20

OpenStudy (muscrat123):

thanks guys

hero (hero):

Great job @musicrat and @Compassionate

OpenStudy (compassionate):

That's correct @muscrat123 ! Thanks, @Hero

OpenStudy (muscrat123):

thanks i cant believe i missed that